OpenStudy (anonymous):

solve for x: 3x + 2(x-4)= 4x - 14 - 2x

OpenStudy (anonymous):

3x+2(x-4)=4x-14-2x 3x+2x-8=2x-14 5x-8=2x-14 3x=-6 x=-2

OpenStudy (anonymous):

that's what I thaught....thx